OpenStudy (anonymous):

A manufacturer produces bolts of fabric with a fixed width. The quantity q of this fabric (measured in yards) that is sold is a function of the selling price p (in dollars per yard), so we can write q = f(p). Then the total revenue earned with selling price p is R(p) = pf(p). Given f(15) = 7000 and f '(15) = −330, find R'(15).

OpenStudy (anonymous):

R'(p) = p f'(p) + f(p) R'(15) = 15 ( -330 ) + 7000 = -4950+ 7000 = 2050